When two electric resistors R1 and R2 are in series, their total resistance (the sum) is 32 ohms. If the same resistors are in parallel, their total resistance (the reciprocal of which equals the sum of the reciprocals of the individual resistances) is the maximum possible for two such resistors. What is the resistance of each?

Answer by rfer(16322)

R1+R2=32
R1=16
R2=16
(R1xR2)/(R1+R2)= 8ohms